OpenStudy (anonymous):

If f(x)=3x-3 and g(x)= x^3 + 27, what is g(f(x))? a. 27x^3 b. 27x^3-81x^2+81x c. 9x^2-18x+36 d. 3x^3+78

OpenStudy (anonymous):

C.

OpenStudy (anonymous):

Wrong answer :(

OpenStudy (anonymous):

It says it was B...

OpenStudy (anonymous):

Dangit.. I guess I'll have to stop going C for cats.. lol.

OpenStudy (anonymous):

Lol

OpenStudy (amistre64):

since g(x) = x^3 + 27 , for any value of x, lets say x=k g(k) = k^3 + 27 if we let k = f(x) = 3x-3 g(k) = k^3 + 27 g(f(x)) = (f(x))^3 + 27 g(f(x)) = (3x-3)^3 + 27 and sinplify